Problem 8-13

Shoney Video Concepts produces a line of video streaming servers that are linked to personal computers for storing movies. These devices have very fast access and large storage capacity.

    Shoney is trying to determine a production plan for the next 12 months. The main criterion for this plan is that the employment level is to be held constant over the period. Shoney is continuing in its R&D efforts to develop new applications and prefers not to cause any adverse feelings with the local workforce. For the same reason, all employees should put in full workweeks, even if that is not the lowest-cost alternative. The forecast for the next 12 months is

MONTH FORECAST DEMAND MONTH FORECAST DEMAND
January 600 July 200
February 800 August 200
March 900 September 300
April 600 October 700
May 400 November 800
June 300 December 900

     Manufacturing cost is $200 per server, equally divided between materials and labor. Inventory storage cost is $5 per month. A shortage of servers results in lost sales and is estimated to cost an overall $20 per unit short.

    The inventory on hand at the beginning of the planning period is 200 units. Ten labor hours are required per server. The workday is eight hours.

Develop an aggregate production schedule for the year using a constant workforce. For simplicity, assume 22 working days each month except July, when the plant closes down for three weeks’ vacation (leaving seven working days). Assume that total annual production capacity is greater than or equal to total annual demand (i.e., compute workforce level based on annual demand and annual capacity). (Leave no cells blank - be certain to enter "0" wherever required. Indicate monthly shortages using a negative ending inventory level. Round up the "number of workers" to the next whole number and round down your "monthly production rates" to the nearest whole number.)


January February March April May June July August September October November December Total
Forecast 600 800 900 600 400 300 200 200 300 700 800 900
Beginning inventory
Available production
Ending inventory
Costs Total
Lost sales
Inventory
Total
Answer #1

Calculations:

F G 1 Beginning inventory 200 2 Hours per day 3 No. of days in July 4 No. of days in other months 22 5 Unit carrying cost 5 6
H M Jul Oct 17 18 200 19 =MAX (0,621) 20 =$B$14 21 =H19+H20-H18 Total =SUM (B18:M18) Aug 200 =MAX(0, H21) =$B$15 =I19+120-118

Results:

Aug May 400 Total 6700 Jan Forecast 600 Beginning inventory200 Available production 580 Ending inventory 180 Feb 800 180 580

----